OpenStudy (yagirlselena):

Solve for Y 8x+2y=24 Need help its for a test tm and idk how to do this

OpenStudy (kailey7277):

\[8x+2y=24\] \[-8x.....-8x\] \[2y=-8x+24\] \[\div2....\div2..\div2\] y=-4x+12

Awolflover1 (awolflover1):

let's go step by step with the user so they can understand.

OpenStudy (yagirlselena):

I'm confused kaily seems righg

OpenStudy (mathmale):

Please remember, helpers, that it is your job to provide guidance so that the other person can solve the problem himself or herself. Do NOT just solve the problem and hand over the answer! That's not teaching.

OpenStudy (mathmale):

First of all, what does "solve for y" mean? Seriously. Explain this in your own words.

OpenStudy (kailey7277):

You have to isolate your variable; you need to get y on one side alone.

OpenStudy (yagirlselena):

I have before she responded 8x+2y=24 +8x +8x 2y = 24+8x

OpenStudy (kailey7277):

So you would first need to subtract 8x from both sides which would give you \[2y=24-8x\]

OpenStudy (yagirlselena):

Yes I had that I just get very confused on what to do next

OpenStudy (kailey7277):

you wouldn't add 8x because if you did that, you would have 16x+29=24+8x

OpenStudy (kailey7277):

After you subtract 8x, you would divide both sides by 2. That way, you would have y alone on one side. When you divide the right side by 2, you need to divide both of the terms by 2.

OpenStudy (mathmale):

Selena: thank you for showing your work. Why did you choose to ADD 8x to both sides? Try subtracting 8x from both sides, and show your work again.

OpenStudy (yagirlselena):

Yes I had add but I meant subtract on my paper was subtract but I was looking at the positive 8x so when I write it I write it +8x+2y=24 -8x =-8x +2y = 24-8x

OpenStudy (kailey7277):

that's correct. now you would divide everything by 2

OpenStudy (mathmale):

Making great progress! Selena, what's your final answer?

OpenStudy (yagirlselena):

Y=12-4x

OpenStudy (mathmale):

Yes, that's right. Practice, practice! Do similar poblems until you feel confident about your ability to do them.